OpenStudy (anonymous):

write and equation of the line passing through each pair of points. (-8,-5), (-3,10)

OpenStudy (anonymous):

Well firstly you have to find the slope which there is a formula for that(y2-y1 / x2-x1) , so it would be (-8,-5) (-3,10) 10-(-5) / -3-(-8) = 5/5= 1 So then you take the formula y-y1=m(x-x1) and plug in whatever pair of coordinates like so : y-(-3)=1(x-10) First you distribute Y-(-3)=1x-10 The double negatives cancel making 3 positive So y+3=1x-10 Add 3 to each side Y= 1x-13

OpenStudy (anonymous):

Oops read a lil wrong , u can do the same thing with the other pair of coordinates